Astronomy Report – February 17th

Crew Astronomer: Ghanim Alotaibi


Robotic Telescope Requested: MDRS-14

Objects Viewed: Last night 10 images were captured out of the requested 18 images for MDRS Target 1. The quality of the images is bad and one of the images looks shaky. Crew astronomer suggested to resubmit the observation and Observatory Director recommended the same due to some problems encountered, thus the observation was resubmitted.

New photometry measurement was performed for SY MON target. Based on the previous experience gained, the new measurement was found to be 14.132 with an error of 0.0185. The data point was submitted to the AAVSO website with the star 122 as a check star.

Problems Encountered: Mount Problems which caused shaky images. The moon was also bright, so bad images were captured. Observatory Director solved the problem.

Copyright © The Mars Society. All rights reserved. | Main Site